Gaseous methane CH4 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 0.160g of methane is mixed with 0.44g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Round your answer to 2 significant digits.


1
Expert's answer
2021-11-17T16:54:02-0500

CH4 + 2O2 → CO2 + 2H2O

M(CH4) = 16.04 g/mol

n(CH4) "= \frac{0.160}{16.0}=0.01 \;mol"

M(O2) = 32.0 g/mol

n(O2) "= \frac{0.44}{32.0}=0.013 \;mol"

For each mole of CH4 we need 2 moles of O2, but for 0.01 mol of CH4 we have only 0.013 mol of O2. So, O2 is the limiting reactant.

According to the reaction:

n(CO2) = "\frac{1}{2}" n(O2) "= \frac{1}{2} \times 0.013 = 0.0065 \;mol"

M(CO2) = 44.0 g/mol

m(CO2) "= 44.0 \times 0.0065 = 0.286 \;g"

Answer: 0.28 g.